OpenStudy (anonymous):

Which fraction is equivalent to 72% ?

OpenStudy (anonymous):

@beccaboo333

OpenStudy (anonymous):

72/100 = 72% 36/50 = 72% 18/25 = 72% 72% = 72/100 or 18/25 in fraction

OpenStudy (anonymous):

thank u!